OpenStudy (anonymous):

steps to solving the solutions of an equation: 6 ( -x + 5) = 2 (x + 8) -10

OpenStudy (alienium):

-6x+30 = 2x + 16 - 10 -6x+30=2x+6 -6x-2x=6-30 -8x=-24 x=3